# Does the square-root price impact law belong to the strict universal scalings?: quantitative support by a complete survey of the Tokyo stock exchange market

### v1

@@ Abstract
Figure 4: Quantitative test of the two theoretical models that have supported the nonuniversality hypothesis: Fig. (a) for the GGPS prediction $\delta=\beta-1$ and Fig. (b) for the FGLW prediction $\delta=\beta-1$. No correlation was found and, thus, both models were rejected.
While we presented direct evidence supporting the universality hypothesis, we now present direct negative evidence against two prominent models supporting the nonuniversality hypothesis: the GGPS and the FGLW models NatureGabaix ; Gabaix2006 ; FGLWQFin2013 .
The GGPS and FGLW models are based on microeconomic theory in the presence of order-splitters: Let us consider the case where an order-splitter has a metaorder with total volume $Q$ and splits it into $L$ child orders. Empirically, the distributions of the metaorder volume $Q$ and the number of corresponding child orders $L$ are known to obey power laws Bershova ; SatoPRL2023 ; SatoPRR2023 ; LMF_PRE2005 ; NatureGabaix ; Gabaix2006 : $P(Q)\propto Q^{-\beta-1}$ for large $Q$ and $P(L)\propto L^{-\alpha-1}$ for large $L$, where $\beta$ and $\alpha$ are positive power-law exponents characterising individual stocks. Based on this key assumption, the GGPS and FGLW models consider the strategy optimisation of market makers in light of inventory risk and fair pricing, respectively. They predict that the exponent $\delta$ is given by
[…]
@@
Interestingly, the typical values of $\beta$ and $\alpha$ are distributed around 1.5, which is superficially consistent with the SRL SatoPRL2023 ; SatoPRR2023 ; NatureGabaix ; Gabaix2006 ; Bershova . However, since $\beta$ and $\alpha$ can differ from 1.5 depending on the stock, the GGPS and FGLW models have suggested the theoretical nonuniversality of the exponent $\delta$. Their prediction (6) can be tested in principle if we can precisely estimate $\beta$, $\alpha$, and $\delta$ with a precision of $0.1$. While a small group of researchers tested this relation (6), both positive and negative results MastromatteoPRE2014 ; NatureGabaix ; Gabaix2006 ; Bershova coexist, as direct precise measurement without bias is not an easy task for $\beta$, $\alpha$, and $\delta$, except in a complete survey. Evidently, concluding this debate has been very challenging without large comprehensive microscopic datasets. In this article, we address this issue definitively using our large TSE dataset.
@@
We investigated the large microscopic dataset on the Tokyo Stock Exchange over eight years and scrutinised the long-standing universality vs. nonuniversality debate regarding the square-root price impact law. We present that (i) the square-root law holds with an exact exponent $\delta=1/2$ within statistical errors both at the stock and trader levels, and (ii) the GGPS and FGLW models – two prominent models theoretically supporting the nonuniversality hypothesis – were rejected. These results provide the strongest evidence so far supporting the universality hypothesis, at least for the Tokyo Stock Exchange. We recommend that interested readers refer to the SI, where we thoroughly checked the statistical robustness of our findings.
Technically, our simulation shows a small bias in our current $\delta$ estimator, such that $e_{\delta}:=\langle\!\langle\delta\rangle\!\rangle-1/2=-0.011$, due to the finite sample-size effect. Given this bias, one of the simplest methods for constructing unbiased estimates of $\delta$’s mean is to use $\langle\delta_{\rm unbiased}\rangle:=\langle\delta\rangle-e_{\delta}$ for bias removal. Consequently, we obtain the unbiased mean
$\langle\delta_{\rm unbiased}\rangle=0.500\pm 0.0020$ (7)
for our dataset, precisely supporting the SRL. However, our numerical null model is too simple to capture the complete price-impact dynamics. Considering the potential modelling error, there may still be room to improve the $\delta$ estimator for faster convergence with exact unbiasedness. This topic is left for future studies.
Our finding implies that microscopic theories behind the square-root law should yield the universal power-law exponent $\delta=1/2$, which does not depend on the details of the financial market microstructure, such as the latent order-book theory BouchaudText ; Toth2011 ; DonierLOB2015 . It would be a future issue to directly validate or falsify such theories using our microscopic dataset.
Additionally, as $\delta$ is universally $1/2$, it would be very interesting to study the system-specific details of the coefficient $c$. Regarding $c$, several natural questions arise: Is it possible to predict $c$ from microscopic financial models? For which types of stocks does $c$ tend to be larger or smaller? If there is a sudden change in the market microstructure, such as a tick-size change, does $c$ change significantly? These questions are very important not only from an academic perspective but also from the viewpoint of practitioners. Indeed, $c$ characterises the true transaction cost associated with large metaorder splitting by institutional investors. Since typical investors aim to minimise transaction costs, the microscopic mechanism behind the square-root law is an exciting topic for practitioners.
While these questions are evidently important, the lack of large microscopic datasets has hindered solid data analyses regarding the coefficient $c$. Indeed, in most of the previous studies on the square-root law, transaction datasets from different stocks were mixed when calculating the average price impact, based on the strong assumption that the dimensional analysis in Eq. (1) is perfectly correct, allowing different datasets to be mixed after non-dimensionalisation. As a side effect, it has been very challenging to obtain the coefficient $c$ for individual stocks. In contrast, since we precisely measured the coefficient $c$ for all stocks individually in this paper, it is now possible, in principle, to investigate the stock-by-stock mechanisms that microscopically determine $c$. Thus, our work strongly motivates the opening of a new field of statistical-physics theories focusing on the coefficient $c$.
Economically, the universality of the square-root law suggests that liquidity should be measured by the coefficient $c$ for large institutional investors. In traditional economics, linear-response theory $I(Q)\approx\lambda Q$ has been assumed, and the linear coefficient $\lambda$ – known as Kyle’s lambda, after Kyle’s pioneering work Kyle – has been used as a liquidity measure. While this metric is plausible for small transactions, the square-root law demonstrates that Kyle’s lambda systematically underestimates the actual liquidity for large metaorder splitting because $Q\gg\sqrt{Q}$ for large $Q$. Thus, our study sheds new light on the quantification of market liquidity in the presence of large institutional investors, an urgent topic relevant to policy evaluation in managing exchange markets.
Acknowledgements
@@ Acknowledgements.
We appreciate the careful reading of this manuscript by the staff of JPX Group. Also, we thank Jean-Phillipe Bouchaud for his motivating comments on our conference presentations and this manuscript. YS was supported by JSPS KAKENHI (Grant No. 24KJ1328). KK was supported by JSPS KAKENHI (Grant Nos. 21H01560 and 22H01141) and the JSPS Core-to-Core Program (Grant No. JPJSCCA20200001).
Author contribution
YS developed the program code and conducted the statistical analysis. KK designed and managed the project. Both YS and KK contributed to writing the manuscript and approved its final content.
Conflict of interest

### v2

@@ Introduction.
Figure 5: Quantitative test of (a) the GGPS prediction $\delta=\beta-1$ and for (b) the FGLW prediction $\delta=\alpha-1$. No correlation was found and, thus, both models were rejected.
While we presented direct evidence supporting the universality hypothesis, we now present direct negative evidence against two prominent models supporting the nonuniversality hypothesis: the GGPS and the FGLW models [24, 25, 26].
The GGPS and FGLW models are based on microeconomic theory in the presence of order-splitters: Suppose that an order-splitter has a metaorder with total volume $Q$ and splits it into $L$ child orders. Empirically, the distributions of the metaorder volume $Q$ and the number of corresponding child orders $L$ are known to obey power laws [31, 28, 29, 30, 24, 25]: $P(Q)\propto Q^{-\beta-1}$ for large $Q$ and $P(L)\propto L^{-\alpha-1}$ for large $L$, where $\beta$ and $\alpha$ are positive power-law exponents characterising individual stocks. Based on this key assumption, the GGPS and FGLW models consider the strategy optimisation of market makers in light of inventory risk and fair pricing, respectively. They predict that the exponent $\delta$ is given by
$\mbox{GGPS: }\delta=\beta-1,\>\>\>\mbox{FGLW: }\delta=\alpha-1,$ (6)
@@
Interestingly, the typical values of $\beta$ and $\alpha$ are distributed around 1.5, which is superficially consistent with the SRL [28, 29, 24, 25, 31]. However, since $\beta$ and $\alpha$ can differ from 1.5 depending on the stock, the GGPS and FGLW models have suggested the theoretical nonuniversality of the exponent $\delta$. Their prediction (6) can be tested by precisely estimating $\beta$, $\alpha$, and $\delta$. While a small group of researchers tested this relation (6), both positive and negative results [7, 24, 25, 31] coexist, as direct precise measurement without bias is not an easy task for $\beta$, $\alpha$, and $\delta$, except in a complete survey. We address this issue definitively using our large TSE dataset.
@@
This Letter provides the strongest evidence so far supporting the universality hypothesis, at least on the TSE. See also the SM [9] for thorough statistical robustness of our findings. It would be desirable to perform complete surveys of other markets, which is left as a future study. However, we stress that a single-market survey is sufficient to falsify the universal validity of the prominent non-universal models, as shown in this work.
Economically, the square-root law implies that market liquidity is larger against metaorder splitting than expected by linear models in traditional economics [37]. Additionally, Ref. [36] implies the SRL plays a key role in validating the efficient market hypothesis—one of the most important concepts in financial economics. Thus, our study sheds new light on market liquidity in the presence of large institutional investors.
@@ Acknowledgements.
Acknowledgements.
We appreciate the careful reading of this manuscript by the staff of JPX Group. Also, we thank Jean-Phillipe Bouchaud for his motivating comments on our conference presentations and this manuscript. YS was supported by JSPS KAKENHI (Grant No. 24KJ1328). $K$ was supported by JSPS KAKENHI (Grant Nos. 21H01560, 22H01141, and 25K01450) and the JSPS Core-to-Core Program (Grant No. JPJSCCA20200001).
Data availability.
The data supporting our results were provided by Japan Exchange (JPX) Group, Inc. JPX Group is a third-party commercial company and provided their dataset through a non-disclosure agreement with Kyoto University, strictly for academic purposes. This non-disclosure agreement imposes legal restrictions on data availability, and therefore, we cannot make the data publicly accessible without approval from JPX Group. However, upon reasonable request, $K$ will provide all the original data and programming code, subject to explicit approval by JPX Group.
Code availability.
Our numerical simulation code, used to estimate the finite sample-size effect on the standard deviation of the estimated exponent $\delta$, is available from the following URL: https://gitlab.com/Yuki-Sato_JPN/CodeDisclosure-2024-SRL.
Ethics statement
The TSE dataset does not contain personal information, as defined by the Act on the Protection of Personal Information in Japan, and no ethical issues are involved in this research. JPX Group defines market participants at the corporate level. JPX Group collected trading log data at the corporate account level by making agreements with corporate-level market participants; in other words, personal information linked to individual traders behind corporate accounts was not collected. Additionally, an ethical review is not required for social science research based on such corporate-account-level datasets at Kyoto University.
Author contribution.
YS developed the program code and conducted the statistical analysis. $K$ designed and managed the project. Both YS and $K$ contributed to writing the manuscript and approved its final content.
Conflict of Interest statement.
@@
Appendix A End Matter
Figure 6: Schematic example of the aggregate scaling plots: (a) Average price impact profile for the three specific stocks from Dataset 2: Morita Holdings Corp. (blue, ticker 6455, $s=1$), FDK Corp. (orange, ticker 6955, $s=2$), and Matsuya Co. LTD. (green, ticker 8237, $s=3$), where $s$ represents the stock index. While this data agrees with the square-root law $\mathcal{I}_{s}(\mathcal{Q})=c_{s}\sqrt{\mathcal{Q}}$ on average, the three coefficient $\{c_{s}\}_{s=1,2,3}$ are not identical. (b) This heterogeneity is removed by applying the rescaling $\phi(x)=\mathcal{I}_{s}(\mathcal{Q}(x))/c_{s}$ with $x:=\sqrt{\mathcal{Q}}$. (c) Next, the average $\phi_{\rm avg}(x)$ is obtained across all the stocks. (d) Finally, the averaged profile $\phi_{\rm avg}(x)$ is plotted with bins evenly spaced on a logarithmic scale.
Appendix A: Data description.
We used the JPX order-book dataset on the TSE. This data covers the complete life cycle of all orders, from submissions to transactions/cancellations, for all stocks on the TSE over eight years (from 4th January 2012 to 2nd November 2019).
We divided this dataset into two datasets: one from 4 January 2012 to 18 September 2015 (Dataset 1), and another from 24 September 2015 to 2 November 2019 (Dataset 2). This is because there was a technical update on the arrowhead—the order-matching system managed by the TSE—on 24 September 2015, and all the virtual server IDs were shuffled.
A liquid stock is defined as one where the total number of metaorders exceeds $10^{5}$. This filter is introduced to control the sample size and the resulting accuracy of estimation. In addition, any stock in Dataset 1 is regarded as different from stocks in Dataset 2, even if their ticker codes are identical. With this definition, the total number of liquid stocks in Dataset 1 and Dataset 2 are 942 and 1,357, respectively. Thus, we have 2,299 datapoints in total regarding liquid stocks.
Our TSE dataset records virtual server IDs corresponding to each order submission. A virtual server acts as an effective unit of trading accounts on the TSE, as all orders are submitted from a virtual server. Technically, a virtual server ID is not exactly technically equivalent to a membership ID, as any membership may possess several virtual servers. In fact, there is a limit on the total number of submissions that can be made during a fixed interval from one virtual server, and high-frequency traders might possess several virtual servers to circumvent this limit. However, we can allocate effective trader IDs by appropriate aggregation of virtual server IDs.
Virtual server ID Order ID Order type Effective trader ID
V1 O1 Limit order T1
V2 O1 Cancellation T1
V3 O2 Limit order T2
V3 O2 Cancellation T2
Table 1: Schematic of how trader IDs are defined from virtual server IDs.
Effective trader IDs are defined as follows: While the lifecycle of a single order ID is typically managed by a single virtual server, there are exceptional cases where multiple virtual servers are involved in the lifecycle of one order. Suppose that virtual server “V1” submits a limit order, and then virtual server “V2” cancels the same limit order under the same order ID “O1” (see Table 1). We infer that the same trader operates both virtual servers, V1 and V2, and we assign a unified identifier, “T1,” as the effective trader ID. This effective trader ID is referred to as a trader ID for simplicity, while it was originally called trading desks in the literature [27, 28, 29].
We define trader IDs for each stock: Even if trader “T1” made transactions for two different stocks (such as Toyota and NTT), we define two traders, such as “Dataset1-Toyota-T1” and “Dataset1-NTT-T1.”
Appendix B: Definition of metaorders.
Metaorders are defined based on the order-splitting hypothesis: typical large investors split a large metaorder into a long sequence of child orders. For example, trader “T1” may have a buy (sell) metaorder with a volume of $Q=+100$ (Q=100Q=-100) and split it into ten buy (sell) child orders, such as $q_{1}=q_{2}=...=q_{10}=+10$ ($q_{1}=q_{2}=...=q_{10}=-10$), where the sign “++” (“-”) represents a buy (sell) order.
Based on this hypothesis, we assume that any two successive orders belong to the same metaorder if their order signs are identical. For example, if we have a market order sequence $\{q_{k}^{(i)}\}_{k}=\{+10,+10,+1,+5,-3,-2,-1,+100\}$ for trader $i$, the $k$-th metaorder volume $Q_{k}^{(i)}$ and the corresponding number of child orders $L_{k}^{(i)}$ are given by $(Q_{1}^{(i)},L_{1}^{(i)})=(+26,4)$, $(Q_{2}^{(i)},L_{2}^{(i)})=(-6,3)$, and $(Q_{3}^{(i)},L_{3}^{(i)})=(+100,1)$. Additionally, any metaorder is defined on a daily basis: market orders on different days are regarded as belonging to separate metaorders.
With this definition, we determined the metaorder volumes and lengths for each stock, and then measured the power-law exponents $\beta$ and $\alpha$ using Clauset’s algorithm [38]. These exponents were used to draw the scatterplots in Fig. 5.
Appendix C: Exceptional handling.
The TSE has exceptional rules known as the daily price limit and circuit breaker, which are aimed at protecting investors from sudden price jumps. Since large price changes are typically caused by significant news, which is beyond the scope of our study, we excluded such exceptional days from our data analysis.
This exceptional handling is anticipated to mitigate selection bias in measuring price impact for large $Q$. For instance, a trader might pause the execution of a buy metaorder if the price goes up unexpectedly fast [3]. Our approach excludes exceptional trading days with sudden price jumps, during which significant implementation bias could exist.
Appendix D: Stock-level price impact.
For a given stock, we defined the daily volatility $\sigma_{\rm D}$ as the difference between the daily maximum and minimum prices [39] (see SM [9] for a robustness check using half-day intervals instead of full-day intervals). Also, we defined the daily transaction volume as $V_{\rm D}$. Using these characteristic daily parameters, we defined the dimensionless volume and corresponding price impact as
$\mathcal{Q}_{k}:=\frac{Q_{k}}{V_{\rm D}},\>\>\>\mathcal{I}_{k}:=\frac{I_{k}}{\sigma_{\rm D}}$ (7)
with the $k$-th metaorder volume $Q_{k}$ and the corresponding price impact $I_{k}$. We then studied the average price impact to apply the nonlinear least-square fitting, such that
$\mathcal{I}(Q):=\langle\mathcal{I}_{k}\mid\mathcal{Q}_{k}=\mathcal{Q}\rangle\approx c\mathcal{Q}^{\delta}$ (8)
with the fitting parameters $c$ and $\delta$ at the stock level. Our analysis focuses exclusively on metaorder splitting through market orders. Including metaorder splitting via limit orders would introduce estimation bias in the price impact, as the price impact of limit orders is theoretically negative.
The fitting process is given as follows: First, we extracted the metaorders and the resulting price impacts. As a filter, we focused only on metaorders whose liquidation time horizon—defined as the physical time of metaorder completion—exceeds 60 seconds. This 60-second rule is inspired by the previous research proposing the 120-second rule [33]. Second, we introduce logarithmically-evenly spaced bins on the $\mathcal{Q}$ axis and take the conditional averages for all bins. Third, we focus only on the bins with sufficient sample sizes (more than 100) and apply the fitting to the conditional averages using the method of relative least squares (for the technical details, see SM [9]). We thus obtained $\delta$ and $c$ for all liquid stocks.
As a robustness check, we also estimated $\delta$ and $c$ using ordinary least squares directly on the scatter plot (without relying on binned averages), and the results remained nearly identical (see SM [9]). This confirms that our statistical findings are highly robust, irrespective of the measurement method employed.
Appendix E: Trader-level price impact.
The method for calculating the trader-level price impact is essentially identical to that used for the stock-level price impact. One important remark is that we focused only on active traders. An active trader is defined as one who meets two conditions: First, the total number of their metaorders must be not less than $10^{4}$. Second, when applying the nonlinear least square fitting to the empirical conditional averages as $\mathcal{I}^{(i)}(Q)\approx c^{(i)}\mathcal{Q}^{\delta^{(i)}}$, there must be at least 10 bins where the sample size exceeds 100. This filer was introduced to estimate $\delta^{(i)}$ and $c^{(i)}$ with sufficient statistical accuracy. With this definition, the numbers of active traders were 1,293 (415 from Dataset 1 and 878 from Dataset 2).
Appendix F: Aggregate scaling plot.
The aim of the aggregate scaling plot is to visualise a single average scaling master curve across many stocks. While plotting a few scaling curves, as in Fig. 2(a), is the simplest method to qualitatively check the validity of the square-root scaling, it is unrealistic to plot more than two thousand curves. Therefore, we use the aggregate scaling plot as a practical option to show the qualitative validity of the square-root scaling across all stocks.
The generating process of the aggregate scaling plot is as follows: First, we obtained the price-impact plots $\mathcal{I}(\mathcal{Q})$ for all stocks. We visually checked the validity of the square-root law for all stocks. Second, we applied the fitting $\mathcal{I}(\mathcal{Q})\approx c\sqrt{\mathcal{Q}}$ with one parameter $c$ for all stocks and obtained the rescaled profiles $\phi(x):=\mathcal{I}(\mathcal{Q})/c$ with a variable transformation $x:=\sqrt{Q}$ across all stocks. Third, while we had many $\phi(x):=\mathcal{I}(\mathcal{Q})/c$ for all stocks, we calculated their average profile as $\phi_{\rm avg}(x)$. Finally, we plotted $\phi_{\rm avg}(x)$ at the points that are evenly spaced along the logarithmic horizontal axis.
Appendix G: Numerical simulation.
Our nonlinear stochastic model with exact $\delta=1/2$ is based on our real dataset and random shuffling. (i) All metaorder execution schedules (in physical time) are identical to those in our real dataset. (ii) However, the signs of the metaorders are randomly shuffled using our real dataset. (iii) Furthermore, the price impact of each metaorder is modeled as a Gaussian random variable, following the square-root law on average with the exact half exponent $\delta=1/2$. The coefficient $c$ and the daily volatility are identical to those in our dataset, but the randomness arises in our model from the shuffling in process (ii) and the Gaussian random variables in process (iii). Since our model has an exact half exponent $\delta=1/2$, it is feasible to evaluate the estimation error of $\delta$ from our real data attributable to the finite sample-size effect, assuming the exact square-root law.